Determine the equation for the tangent to the curve

??3 - ?3? = 6 at the point (1, 2).
Answer #1

solution is The given curve xy3 – x3y=6 0 Now, differentiating o with rested we get y² + 3xy² dy ca 3x²y_xzdy - o da dy or, dto is Hence the equation for the tangent the given curve at the point (1,2) y-2 =ů (2-3 ing 9-2=-2X + 2 + î + 2 2 11 ie y = 1